The problem describes a debt to be amortized. (Round your answers to the nearest cent.)

Sean Lee purchases $40,000 worth of supplies for his restaurant by making a $3,000 down payment and amortizing the remaining cost with quarterly payments over the next 4 years. The interest rate on the debt is 16% compounded quarterly.

(a)

Find the size of each payment.

(b)

Find the total amount paid for the purchase

(c)

Find the total interest paid over the life of the loan.
Answer #1

Cost = $40000 Down Payment = $3000 .. Loan = 40000 - 3000 37000 = Time = Rate = 4 16Y. years, compounded quarterly- P= r (PU) (trgon P= payment r= rate per period = 16% = na no. of periods = 4x4 = 16 0.04 .. pa oon [ 37000] [lto.04] 1480 0